A ball is dropped from a balloon going up at a speed of 10m/s .If the balloon was at a height 75m at the time of dropping the ball ,how much time will the ball take in reaching the ground ?

Grade:

2 Answers

Arun
4 years ago
Initially ball is going upwards 
Initial velocity=u= -10m/s
S=75m
a=g=10m/s2
From Second equation of motion 
s=ut+1/2at²
75= - 10t +(1/2 )10 t²
75=-10t +5t²
5t² -10t -75=0
t= 5 seconds
∴ The ball will take 5 sec to reach the ground.
 
Khimraj
4 years ago
Solution :
Initially ball is going upwards 
Initial velocity=u= -7m/s
S=75m
a=g=10m/s2
From Second equation of motion 
s=ut+1/2at²
70= - 7t +(1/2 )10 t²
70=-7t +5t²
5t² -7t -70=0
t=-7±√49-4x5(-70)/2x5
t=-7±√49+1400/10
t =5 sec [ as t cannot be negative]
∴ The ball will take 5 sec to reach the ground.